# Vectorized Computation of Euler Characteristic Functions and Transforms

### v1

@@ 2.3 Existing Frameworks
GPU ECC [wang2022gpu] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems.
@@
In this paper, we present the pyECT framework for computing the WECT. In general, the pyECT framework has several advantages over these existing implementations. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across axis-aligned cubical complexes, arbitrary cubical complexes, and geometric simplicial complexes; see Table 1. Second, pyECT provides frameworks for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the vectorized operations used in pyECT are highly optimized for modern GPU hardware, which, in most cases, results in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the simplicity of the vectorized approach makes it very advantageous for software implementation, especially in the context of well-maintained, abstracted tensor computation libraries like PyTorch [pytorch1, ansel2024pytorch]. For example, the core WECT algorithm in the pyECT implementation is only about $25$ lines of code.
@@
One minor disadvantage of pyECT is that it does not compute exact Euler characteristic values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT because directional WECFs do not have consistent heights at which they change values across varying directions. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [Jiang2020TheWE, Turner2014PHT, munch2023invitation, amezquita2022measuring]). While it is possible to compute a vectorization from exact values using EuCalc, it requires an additional computation step that is linear in the size of the discretization.
@@
Package
Simplicial
Complexes
Cubical
Complexes
Axis-Aligned
Cubical Complexes
@@ 2.4 Tensor Operations
Our algorithms almost exclusively use vectorized operations on tensors. These have two main advantages over non-vectorized approaches. The first advantage is that the vectorized operations we use have highly optimized implementations in PyTorch [pytorch1, ansel2024pytorch] that run in parallel on both CPUs and GPUs. The second advantage is that they support automatic gradient computations, facilitating the use of our implementation in neural networks. Here, we review tensors and the vectorized operations used in our algorithm.
@@
We exclusively use one-dimensional tensors (i.e., vectors), two-dimensional tensors (i.e., matrices), and three-dimensional tensors in our algorithm. A specific dimension of a tensor is called an axis.
Next, we introduce the tensor operations utilized in our algorithm. For examples of the following tensor operations, we refer the reader to Appendix A.
@@
    In this work, we only use the cumulative sum on two-dimensional tensors over the second axis, $m_{1}]]$, as defined here. In more general settings, the cumulative sum can be defined over any axis of a tensor.
@@
    Reduce maximum: Let $T$ be a tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ and let $p\in[[d]]$. Then, $\texttt{ReduceMax}(T,p)$ is the tensor with shape
    $(m_{0},\ldots,m_{p-1},m_{p+1},\ldots,m_{d-1})$
    and values
@@
    That is, for a tensor $S$, we have
    $\texttt{ScatterAdd}(I,T)(S)=S+D,$
    where ++ is taken element-wise and $S$ is updated in place. In practice, the tensor $D$ is not explicitly built, but is helpful for analysis.
@@ 3 Computation of WECFs
In this section, we present a new algorithm for computing the WECFs of a weighted simplicial complex with respect to a finite set of vertex filters. Computation of the WECT of a weighted geometric simplicial complex is the motivation for Algorithm 1, and corresponds to the case where the vertex filters are height functions. See Example 3.1 for specifics on how Algorithm 1 may be used to compute the WECT.
3.1 Algorithm
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m}\}$ for the entirety of this section. For any $i\geq 0$, we denote by $K_{i}$, the set of $i$-simplices in $K$ and we let $k_{i}=|K_{i}|$. We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}}^{i}\}$ for each $i\geq 1$. To simplify notation, we extend the vertex filters $f_{p}:K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}:K\to\mathbb{R}$ as
@@ 3.1 Algorithm
        Filters is a tensor with shape $(k_{0},m)$ and values
        $\texttt{Filters}[a,p]=f_{p}(v_{a}).$
        VertexWeights is a tensor with shape $(k_{0})$ and values
        $\texttt{VertexWeights}[a]=\omega(v_{a}).$
@@
        $i\texttt{-SimplexWeights}$ is a tensor with shape $(k_{i})$ and values
        $i\texttt{-SimplexWeights}[b]=\omega(\sigma_{b}^{i}).$
@@
Define MaxVal as
@@
The function $\beta$ has a left inverse
$\alpha\colon[-\texttt{MaxVal},\texttt{MaxVal}]\to[[\texttt{NumVals}]]$
defined as
@@
for all $t\in[-\texttt{MaxVal},\texttt{MaxVal}]$ and $q\in[[\texttt{NumVals}]]$. In other words, $\alpha$ and $\beta$ form a Galois connection.
@@
0:WECFs: tensor of shape $(m,\texttt{NumVals})$ and values $\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}(\beta(q))$.
@@
After the conclusion of the for loop, we have
$\texttt{diffWECFs}=\sum_{i=0}^{\dim(K)}D_{i}.$
Finally, a cumulative sum is applied to diffWECFs to obtain the tensor WECFs. From the fact that the cumulative sum is a linear operator, we have
@@
Likewise, for any $i\in\{1,\ldots,\dim(K)\}$, Equation (3.1) together with the same argument used to show Equation (3.1) gives
@@
Example 3.1.
Computing a discretization of the WECT of a weighted geometric simplicial complex $\mathbb{K}=(K,\omega)$} in $\mathbb{R}^{n}$ is a special case of Algorithm 1 as follows. Consider a tensor $D$, with shape $(d,n)$, of direction vectors over which to sample the WECT. For any $p\in[[d]]$, let $\mathbf{s}_{p}\in\mathbb{R}^{n}$ be the $p$-th row of $D$. Let $V$ be the tensor, with shape $(k_{0},n)$, of the vertex coordinates of $K$. Then, the tensor $V*D^{T}$ has values
@@
3.2 Complexity
Here, we analyze the asymptotic complexity of Algorithm 1. We make the assumption that the ambient dimension $n$ is fixed. Let $k$ be the total number of simplices of $K$, let $m$ be the number of vertex filters, and let NumVals be the number of values to sample the vertex filters over.
Next, the operation
$\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$
is applied to diffWECFs. ScatterAdd adds a value from VertexWeights to diffWECFs for each entry of $\texttt{VIndices}^{T}$, resulting in a time complexity of $\Theta(km)$ and no additional memory usage as the operation is performed in place.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
which takes $\Theta(km)$ time and memory.
After the conclusion of the for-loop, we have a total time and memory complexity of
$\Theta\big(nkm\big)=\Theta(km).$
Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in a final time and memory complexity of $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1.
In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a theoretically optimal final time and memory complexity of $\Theta(k)$ for Algorithm 1.
@@ 4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods on a variety of datasets.
@@ 4.1 Experimental Design
The following section outlines the experimental design, including the datasets used, the algorithms compared, and the parameters used in the experiments. Experiments were performed on a single node with a single CPU or a single GPU (where applicable). The CPU used was an EPYC Rome, and the GPU used was an NVIDIA A40. Run times were recorded using the timeit Python library with GPU synchronization (to prevent inaccurate timing from asynchronous GPU operations).
Datasets
The methods were evaluated on both image and three-dimensional point cloud datasets. In particular, we use:
  • Fashion-MNIST [xiaoFashionMNIST]. Fashion-MNIST is a dataset consisting of $28\times 28$ grayscale images of clothing items. In our experiments we use the training set, which consists of $60,000$ images.
  • ImageNet [deng2009ImageNet]. The Large Scale Visual Recognition Challenge (ILSVRC) 2012-2017 subset of ImageNet was used. From the $50,000$ validation images, we randomly selected $1,000$ images with a dimension (horizontal or vertical) of between $200$ and $500$ pixels. This creates a dataset of images that generate complexes large enough to demonstrate the computational efficiency of our method, while still being small enough to avoid memory constraints in our computation environment.
Implementations
Section 2.3 outlines a variety of existing methods for computing topological functions and transforms. As discussed in Remark 2.2, methods like EuCalc, Demeter, and SINATRA compute the directional (W)ECF and (W)ECT, while methods like FastTopology and GPU ECC compute the ECF. For computation of the WECT, we compare our method with EuCalc  [^4] The implementation of EuCalc was accessed at https://github.com/HugoPasse/Eucalc. . Note that Demeter and SINATRA are not optimized for speed, and are thus not included in the comparison. Furthermore, Demeter is included in a comparison with EuCalc in [lebovici2024efficientcomputationtopologicalintegral], where the implementation is shown to be significantly slower than EuCalc. In the case of computing the ECF, we compare our method with FastTopology  [^5] The implementation of FastTopology was accessed at https://github.com/zavalab/ML/tree/master/FastTopology. . Note that the implementation of GPU ECC is only available on Windows operating systems, which was not available in our computation environment.
@@ 4.2 Results
We now present the experimental results, measured by the speedup factor, of our method over the existing methods. Speedup factor is defined as the ratio of the average runtime of the existing baseline method (either EuCalc or FastTopology) divided by the average runtime of the proposed method (pyECT). The first section explores the computation of the WECT, while the second section explores the computation of the ECF.
4.2.1 WECT Computation
In this section, we present the results of computation of the WECT using our vectorized method. We provide a comparison of the runtime of our method with EuCalc on the Fashion-MNIST and ImageNet datasets. While pyECT significantly outperforms EuCalc on the Fashion-MNIST dataset, the speedup is far more pronounced on the ImageNet dataset. This is because Fashion-MNIST consists of $28\times 28$ pixel images while the images we selected from ImageNet are significantly bigger.
The results for Fashion-MNIST are shown in Figure 2. As the number of directions increases, the speedup of our method over EuCalc increases linearly. We note that increasing the number of heights used in the vectorization of our method (which is not a parameter of EuCalc) does not have a significant effect on the speedup. At $25$ directions, our method is approximately twice as fast as EuCalc.
Figure 2: Speedup of pyECT over EuCalc on WECT computation of Fashion-MNIST.
@@ 4.2.1 WECT Computation
The results for ImageNet are shown in Figure 3. As with Fashion-MNIST, as the number of directions increases, the speedup of our method over EuCalc increases. However, as the complexes are significantly larger, for $25$ directions the speedup of our method ranges from $25$ times to $170$ times. As EuCalc does not explicitly depend on the number of heights, its runtime remained constant as the number of heights used increases. However, our method’s speedup factor increases as the number of heights increases (despite requiring more vector computation). While this appears counterintuitive, in reality the vectorized code encounters less efficient memory access patterns at lower numbers of heights, which leads to a smaller advantage over EuCalc with the smaller tensor values. For a technical description of memory access patterns specific to the GPU hardware used in these experiments, we refer the reader to [gpu2025].
Figure 3: Speedup of pyECT over EuCalc on WECT computation of ImageNet.
For both of these datasets, pyECT demonstrates a significant speedup over EuCalc. Note, however, that EuCalc computes exact height values for the WECT, whereas our method computes a vectorization of the WECT. Of course, in many applied settings, vectorizations are more desirable than exact heights due to their utility in machine learning and statistical pipelines. Obtaining a vectorization from EuCalc involves an additional step that further slows it down in comparison to pyECT. The above experiments demonstrate that pyECT admits large speedups when compared to EuCalc, even without the additional step of producing a vectorization in EuCalc.
4.2.2 ECF Computation
In this section, we present the results of computation of the ECF using our vectorized method. We compare our method with FastTopology on the Fashion-MNIST and ImageNet datasets. Additionally, a “padded” ImageNet dataset was added, which increased the dimensions of each image to a size of $1000\times 1000$ pixels, with intensities drawn uniformly at random. The padded dataset was added because for smaller datasets (e.g., Fashion-MNIST), the overhead of parallelization can mitigate the computational speedups due to the smaller size of the complexes. For pyECT, the number of heights was fixed to $256$ to match that of FastTopology (and compute across each intensity). The results are presented in Figure 4.
Across all datasets, pyECT demonstrates a significant speedup over FastTopology. For the single core case, due to the small size of the complexes in Fashion-MNIST, the results of the two methods are comparable. However, in the single core case for the larger datasets, pyECT demonstrates a speedup of nearly $20$ times for the unpadded dataset, and nearly $30$ times for the padded dataset. As the number of cores used in FastTopology increased, the advantage of pyECT decreased slightly. However, pyECT still demonstrated a speedup of nearly $5$ times for Fashion-MNIST, $10$ times for ImageNet, and $20$ times for the padded ImageNet dataset.
Figure 4: Speedup of pyECT over FastTopology across multiple cores of ECF computation on Fashion-MNIST, ImageNet, and a padded ImageNet

### v2

@@ 2.3 Existing Frameworks for Computation of the ECF and (W)ECT
The FastTopology framework [^5] https://github.com/zavalab/ML/tree/master/FastTopology [13] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF. GPU ECC [24] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems (which is unavailable in our computation environment).
@@
The framework presented in this paper, pyECT, admits several advantages over these existing packages for computing the WECT and ECF. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across axis-aligned cubical complexes, arbitrary cubical complexes, and geometric simplicial complexes; see Table 1. Second, pyECT provides frameworks for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the vectorized operations used in pyECT are highly optimized for modern GPU hardware, which, in most cases, results in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the simplicity of the vectorized approach makes it very advantageous for software implementation, especially in the context of well-maintained, abstracted tensor computation libraries like PyTorch [16, 2]. For example, the core WECT algorithm in the pyECT implementation is only about $25$ lines of code.
@@
One minor disadvantage of pyECT is that it does not compute exact EC values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [12, 23, 15, 1]). While it is possible to compute a vectorization from exact values using EuCalc, it requires an additional computation step that is linear in the size of the discretization.
@@
Package Simplicial Complexes Cubical Complexes Axis-Aligned Cubical Complexes
@@ 2.4 Tensor Operations
Our algorithms almost exclusively use vectorized operations on tensors. These have two main advantages over non-vectorized approaches. First, the vectorized operations we use have highly optimized implementations in PyTorch [16, 2] that run in parallel on both CPUs and GPUs. Second, they support automatic gradient computations, facilitating the use of differentiable versions in our implementations in neural networks. Here, we review tensors and the vectorized operations used in our algorithm.
@@
We exclusively use one-dimensional tensors (i.e., vectors), two-dimensional tensors (i.e., matrices), and three-dimensional tensors. A specific dimension of a tensor is called an axis. We now introduce the tensor operations utilized in our algorithm (and refer the reader to Appendix A for examples).
@@
    In this work, we use the cumulative sum on two-dimensional tensors over the second axis, $m_{1}]]$. However, the cumulative sum can be defined over any axis of a tensor.
@@
    Reduce maximum: Let $T$ be a tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ and let $p\in[[d]]$. Then, $\texttt{ReduceMax}(T,p)$ is the tensor with shape $(m_{0},\ldots,m_{p-1},m_{p+1},\ldots,m_{d-1})$ and values
@@
    That is, for a tensor $S$, we have $\texttt{ScatterAdd}(I,T)(S)=S+D,$ where ++ is taken element-wise and $S$ is updated in place. In practice, the tensor $D$ is not explicitly built.
@@ 3 Computation of WECFs
In this section, we present a new algorithm for computing the WECFs of a weighted simplicial complex with respect to a finite set of vertex filters. Computation of the WECT of a weighted geometric simplicial complex is a special case of Algorithm 1, and corresponds to the case where the vertex filters are height functions; see Example 11.
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m-1}\}$ for the entirety of this section. To simplify notation, we extend each $f_{p}:K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}:K\to\mathbb{R}$ as
@@
We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}}^{i}\}$ for each $i\geq 1$.
@@
        Filters is a tensor with shape $(k_{0},m)$ and values $\texttt{Filters}[a,p]=f_{p}(v_{a}).$
        VertexWeights is a tensor with shape $(k_{0})$ and values $\texttt{VertexWeights}[a]=\omega(v_{a}).$
@@
        $i\texttt{-SimplexWeights}$ is a tensor with shape $(k_{i})$ and values $i\texttt{-SimplexWeights}[b]=\omega(\sigma_{b}^{i}).$
@@
In order to discretize the filtration height parameter, define MaxVal as
@@
The function $\beta$ has a left inverse $\alpha\colon[-\texttt{MaxVal},\texttt{MaxVal}]\to[[\texttt{NumVals}]]$ defined as
@@
for all $t\in[-\texttt{MaxVal},\texttt{MaxVal}]$ and $q\in[[\texttt{NumVals}]]$. [^6] In other words, $\alpha$ and $\beta$ form a Galois connection.
@@
0:WECFs: a tensor of shape $(m,\texttt{NumVals})$ and values $\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}(\beta(q))$.
@@
After the conclusion of the for loop, we have $\texttt{diffWECFs}=\sum_{i=0}^{\dim(K)}D_{i}.$ Finally, a cumulative sum is applied to diffWECFs to obtain the tensor WECFs. From the fact that the cumulative sum is a linear operator, we have
@@
Likewise, for any $i\in\{1,\ldots,\dim(K)\}$, Equation (3) together with the same argument used to show Equation (3) gives
@@
Example 11.
Computing a discretization of the WECT of a weighted simplicial complex $\mathbb{K}=(K,\omega)$} with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$ is a special case of Algorithm 1 as follows. Consider a tensor $D$, with shape $(d,n)$, of direction vectors over which to sample the WECT. For any $p\in[[d]]$, let $\mathbf{s}_{p}\in\mathbb{R}^{n}$ be the $p$-th row of $D$. Let $V$ be the tensor, with shape $(k_{0},n)$, of the vertex coordinates of $K$. Then, the tensor $V*D^{T}$ has values
@@
The algorithm achieves a theoretically optimal final time and memory complexity of $\Theta(k)$, where $k$ is the number of simplices in $K$. We refer the reader to Appendix B for the full complexity analysis of Algorithm 1.
@@ 4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods; see Section 2.3 on a variety of datasets. [^7] The implementation of pyECT can be accessed here: https://github.com/compTAG/pyECT. The source code for reproducing experiments (with instructions) is available here: https://github.com/compTAG/pyECT-experiments.
@@ 4.1 Experimental Design
Experiments were performed on a single node with a single CPU or a single GPU (unless otherwise specified). EPYC Rome CPUs and NVIDIA A100 GPUs were used. Runtimes for CPU experiments were recorded using the time Python library using the performance counter for benchmarking. For the GPU experiments, runtimes were recorded using PyTorch’s torch.cuda.Event API to prevent inaccurate timing from asynchronous GPU operations. The methods were evaluated on the following two- and three-dimensional datasets:
  • Fashion-MNIST [25]. Fashion-MNIST is a dataset consisting of $28\times 28$ grayscale images of clothing items. Its training partition consists of $60000$ images with equal representation across ten classes ($6000$ images per class). In our experiments, we use a random subset of $1000$ images sampled uniformly without replacement from the training partition, which are then converted to cubical complexes.
  • ImageNet [7, 19]. The ImageNet Sample Image dataset [19], which consists of $1000$ samples from the Large Scale Visual Recognition Challenge (ILSVRC) 2012-2017 subset of the ImageNet dataset [7] was used. All images in the sample were converted to grayscale using the ITU-R 601-2 luma transform within the Pillow Python library [4]. Then, the images were padded to the maximum dimensions of all images in the dataset, resulting in each image having uniform size of $1632\times 2400$ pixels. The padded pixels were assigned a pixel intensity value of $1.0$ (on a zero to one scale), and the images were converted to cubical complexes.
  • 3DCCs. A synthetic dataset of $1000$ three-dimensional cubical complexes was generated. Each complex has dimension $100\times 100\times 100$, with pixel intensities drawn i.i.d. from the uniform distribution on $[0,1]$.
  • Stanford Bunny and Armadillo [21]. The Stanford 3D Scanning Repository provides triangulated mesh models of three-dimensional objects. We use the Bunny and Armadillo models. The meshes were converted to (unweighted) simplicial complexes after preprocessing to de-duplicate vertices, edges, and faces. In the resulting simplicial complexes, the Bunny dataset consists of $2503$ vertices, $7473$ edges, and $4968$ faces. The Armadillo is significantly larger, consisting of $172974$ vertices, $518916$ edges, and $345944$ faces.
@@ 4.2 Results
We now present the experimental results, measured by the average speedup factor of our method over the existing methods. The average speedup factor is defined as the runtime of the baseline method (e.g., EuCalc, DECT, or FastTopology) divided by the runtime of the proposed method, pyECT, averaged across all images in the dataset.
WECT Results
In this section, we present results on the average speedup of pyECT over EuCalc and DECT when computing the WECT of the Fashion-MNIST and ImageNet datasets. As EuCalc does not provide a GPU implementation, all results for EuCalc are measured on the CPU. Also, as EuCalc does not natively provide a vectorized result, we provide results both including and excluding the vectorization time.
The results for Fashion-MNIST are shown in Figure 2. Across all settings, pyECT demonstrates speedups over both EuCalc and DECT. When comparing against EuCalc (Figure 2(2(a))–(2(d))), pyECT demonstrates a speedup in both CPU and GPU implementations—even without the vectorization step. However, the GPU implementation of pyECT significantly outperforms the CPU implementation as the number of directions increases, which is especially pronounced when including the vectorization time for EuCalc. As expected, when increasing the number of values used in the filtration, the speedup factor of pyECT over EuCalc increases drastically. When comparing against DECT (subfigures (2(e)) and (2(f))), pyECT also demonstrates significant speedups. While the computation speed is comparable for low numbers of directions and filtration heights, pyECT significantly outperforms DECT for larger choices of parameters. Also, as DECT is optimized for GPU architectures, the average speedup on GPU architectures is less pronounced than that of other comparisons. However, when constrained to CPU architectures, pyECT is much faster, indicating another advantage of pyECT’s general applicability for all scenarios.
(a) pyECT (CPU) vs. EuCalc
(b) pyECT (GPU) vs. EuCalc
(c) pyECT (CPU) vs. EuCalc (vectorized)
(d) pyECT (GPU) vs. EuCalc (vectorized)
(e) pyECT (CPU) vs. DECT (CPU)
(f) pyECT (GPU) vs. DECT (GPU)
Figure 2: A comparison for computing on the Fashion-MNIST dataset for WECT computation. Figures (2(a)) and (2(b)) compare pyECT with EuCalc excluding the vectorization time, (2(c)) and (2(d)) compare pyECT with EuCalc including the vectorization time, and (2(e)) and (2(f)) compare pyECT with DECT. CPU implementations of pyECT are shown on the left and GPU implementations are shown on the right. Observe that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights.
@@
The average speedup results on the ImageNet are shown in Figure 3. When comparing against EuCalc (both with and without the vectorization step, in Figures 3(a)3(d)), pyECT demonstrates a significant speedup in both CPU and GPU implementations. This speedup is especially pronounced for the GPU implementation of pyECT, as it is much more stable to large numbers of directions and filtration height values. Due to the large size of the complexes in this dataset, the CPU speedup of pyECT is less pronounced than on the Fashion-MNIST dataset (especially for large numbers of directions and heights). Furthermore, the vectorization time of EuCalc is less impactful on this dataset, as the vectorization step is less significant compared to the overall runtime for the large complexes. In the GPU implementation of pyECT, we observe a trend where the average speedup increases as the number of height filtration values increases. While this behavior appears counterintuitive, in reality the vectorized code encounters less efficient memory access patterns at lower numbers of heights, which leads to a smaller advantage over EuCalc with the smaller tensor values. For a technical description of memory access patterns specific to the GPU hardware used in these experiments, we refer the reader to [8]. In the comparison against DECT (Figures 3(e) and 3(f)), pyECT again demonstrates significant speedups. In fact, due to the large size of the complexes in this dataset, DECT was unable to compute the WECT for large numbers of directions and height values (due to memory constraints on the GPU). This results in the appearance of missing values in Figures 3(e) and 3(f). Note, however, that pyECT is still able to compute the WECT in these scenarios, demonstrating its memory efficiency. Furthermore, in the cases where the WECT could be computed by DECT, the pyECT implementation outperforms DECT for larger numbers of directions and height filtration values.
(a) pyECT (CPU) vs. EuCalc
(b) pyECT (GPU) vs. EuCalc
(c) pyECT (CPU) vs. EuCalc (vectorized)
(d) pyECT (GPU) vs. EuCalc (vectorized)
(e) pyECT (CPU) vs. DECT (CPU)
(f) pyECT (GPU) vs. DECT (GPU)
Figure 3: A comparison for computing on the ImageNet dataset for WECT computation. Figures (3(a)) and (3(b)) compare pyECT with EuCalc excluding the vectorization time, (3(c)) and (3(d)) compare pyECT with EuCalc including the vectorization time, and (3(e)) and (3(f)) compare pyECT with DECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights. Note that DECT was unable to compute the WECT for large numbers of directions and heights due to memory constraints, resulting in missing values in (3(e)) and (3(f)). However, pyECT was able to compute the WECT in these scenarios.
Speedup results for the comparison of pyECT against DECT on the three-dimensional mesh datasets are presented in Figure 4. Note that experiments were run 10 times on each mesh dataset to generate the average and standard deviation values shown in the plots. For large choices of directions and filtration heights in both datasets, DECT was unable to compute the ECT due to memory constraints. While this results in missing values in all subplots, from the cases where DECT was able to compute the ECT, pyECT outperforms DECT (especially for larger choices of parameters). The memory issues are more pronounced in the Armadillo dataset (due to its larger size). Note that while pyECT demonstrates significant speedups in the GPU implementations, its advantage is even more pronounced in the CPU implementations, highlighting the versatility of pyECT across different hardware architectures.
(a) Bunny: pyECT (CPU) vs. DECT (CPU)
(b) Bunny: pyECT (GPU) vs. DECT (GPU)
(c) Armadillo: pyECT (CPU) vs. DECT (CPU)
(d) Armadillo: pyECT (GPU) vs. DECT (GPU)
Figure 4: A comparison of pyECT against DECT for computing the ECT on the Stanford Bunny and Armadillo meshes. Note that vertical axes are not consistent across subfigures. In many cases, DECT was unable to compute the ECT due to memory constraints (leading to missing values in the plots). However, pyECT was able to compute the ECT in all scenarios, and especially outperforms DECT for large numbers of directions and filtration heights.
ECF Results
In this section, we present the speedup of pyECT over FastTopology when computing the ECF of the Fashion-MNIST, ImageNet, and 3DCCs datasets. Due to the consistent computation graph structure of the ECF algorithm, pyECT is able leverage the PyTorch TorchDynamo compilation framework. These compiled runtimes are also measured and compared with FastTopology. As FastTopology is parallelizable across multiple CPU cores, we include results for FastTopology using $1$, 22, 44, 88, $12$, and $16$ CPU cores. Note that FastTopology does not provide a GPU implementation. The results are presented in Figure 5. We observe that pyECT provides a significant speedup over FastTopology in nearly all cases. When constrained to CPU architectures, pyECT has a less pronounced speedup, but is still comparable in speed on a single CPU core when compared with the fully parallelized FastTopology implementation. When utilizing GPU architectures, pyECT demonstrates a significant speedup over the fully parallelized FastTopology, especially on the three-dimensional 3DCCs dataset. The compilation strategy further increases the speed of pyECT on GPU architectures, but adds unrewarded overhead on CPU architectures.
(a) pyECT (CPU) vs. FastTopology
(b) pyECT (GPU) vs. FastTopology
(c) pyECT (compiled, CPU) vs. FastTopology
(d) pyECT (compiled, GPU) vs. FastTopology
Figure 5: A comparison of average speedup of pyECT over FastTopology when computing the ECF of multiple datasets. The comparison evaluates both uncompiled and compiled versions of pyECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms FastTopology. This behavior is especially pronounced when using the GPU architecture on the three-dimensional 3DCCs dataset.